Question 19 Given f(x) = x2 +2 and g(x) = 3x - 4 find (fog)(x) 9x2 - 12x+9 3x3-8 09x2 - 24x + 18 O x² + 3x-2
Answer #1

19) frx) = x+2 91x1= 3X-4 %3D foglal= f{)5 -(3x-4)+2 9x+16-24u+2 9x-24x+18 Ars %3D
Answer #1

19) frx) = x+2 91x1= 3X-4 %3D foglal= f{)5 -(3x-4)+2 9x+16-24u+2 9x-24x+18 Ars %3D
Answer #1

19) frx) = x+2 91x1= 3X-4 %3D foglal= f{)5 -(3x-4)+2 9x+16-24u+2 9x-24x+18 Ars %3D